# MedArabiQ: Benchmarking Large Language Models on Arabic Medical Tasks

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated significant promise for various applications in healthcare. However, their efficacy in the Arabic medical domain remains unexplored due to the lack of high-quality domain-specific datasets and benchmarks. This study introduces MedArabiQ, a novel benchmark dataset consisting of seven Arabic medical tasks, covering multiple specialties and including multiple choice questions, fill-in-the-blank, and patient-doctor question answering. We first constructed the dataset using past medical exams and publicly available datasets. We then introduced different modifications to evaluate various LLM capabilities, including bias mitigation. We conducted an extensive evaluation with five state-of-the-art open-source and proprietary LLMs, including GPT-4o, Claude 3.5-Sonnet, and Gemini 1.5. Our findings highlight the need for the creation of new high-quality benchmarks that span different languages to ensure fair deployment and scalability of LLMs in healthcare. By establishing this benchmark and releasing the dataset, we provide a foundation for future research aimed at evaluating and enhancing the multilingual capabilities of LLMs for the equitable use of generative AI in healthcare.
Data and Code Availability
In this article, we present a new benchmark dataset, MedArabiQ, and evaluate state-of-the-art LLMs. We release our data for reproducibility, to enable a fair evaluation of LLM models in the future: https://github.com/nyuad-cai/MedArabiQ
Institutional Review Board (IRB)
This study does not involve human subjects, so IRB approval was not required.
@@ 1 Introduction
The use of LLMs in healthcare has sparked enthusiasm due to the potential to improve diagnostic processes, clinical decision-making, and overall patient care (Nazi and Peng, 2023; Meng et al., 2024). One particular application of interest is medical education, in which LLMs can generate concise summaries of information and support interactive learning experiences (Benítez et al., 2023). To this end, several benchmarks have been proposed to assess the capabilities of LLMs in medical knowledge and reasoning. Despite these advances, challenges remain, including ethical concerns, risks of generating biased or harmful content, and variability in performance in different languages and cultural contexts (Yang et al., 2024b; Nazi and Peng, 2023).
@@
Existing benchmarks, such as GLUE and MedQA, cater primarily to English, leaving a significant gap in the evaluation of LLMs for Arabic healthcare applications (Qiu et al., 2024). This is due to multiple factors including the limited availability of high-quality Arabic datasets for clinical applications, coupled with the unique linguistic complexity of Arabic, especially when considering the many different Arabic dialect regions (e.g., Gulf, Maghreb, Egypt, Levant, among others) in addition to Modern Standard Arabic (Salameh et al., 2018). Additionally, although various multilingual models include Arabic in their training data, their performance often falls short in clinical contexts due to insufficient domain-specific resources and a lack of appropriate benchmarks (Nazi and Peng, 2023; Gangavarapu, 2024). Addressing these gaps is critical to unlocking the full potential of LLMs for Arabic-speaking patients and providers, ensuring equitable access to AI-driven advancements in healthcare.
@@
To address these gaps, there is a growing need for frameworks that assess LLM performance in clinical tasks specific to Arabic-speaking populations. By developing benchmarks that reflect real-world clinical interactions, we can ensure more reliable and culturally appropriate LLM deployment in multilingual healthcare systems. In this study, we make several key contributions to address these challenges by introducing MedArabiQ (see Figure 1). First, we developed seven benchmark datasets designed to evaluate LLMs in Arabic healthcare applications while addressing Arabic linguistic complexity along with the domain-specific challenges hindering existing models. We focused on critical healthcare medical tasks, including medical question answering, clinical dialogue, and ethical decision-making. Second, we analyzed the performance of both multilingual and Arabic LLMs, highlighting the impact of linguistic coverage and training data transparency on healthcare applications. We conducted comprehensive evaluation to assess model performance, in order to provide a robust foundation for advancing AI-driven solutions for Arabic healthcare tasks.
Figure 1: Overview of MedArabiQ. We constructed seven new benchmark datasets and evaluated five state-of-the-art LLMs on them.
@@ 2 Related Work
Several benchmarks have been proposed for the evaluation of LLMs for various medical tasks, predominantly in the English language. For example, Gao et al. (2023) introduce Dr. Bench, a diagnostic reasoning benchmark for clinical NLP emphasizing clinical text understanding, medical knowledge reasoning, and diagnosis generation. The benchmark combines English-language data from various sources but primarily consists of in-hospital clinical notes. To assess LLM performance in medical question-answering tasks, a number of benchmarks incorporated samples from medical board exams, such as MedQA. MedQA advances multilingual benchmarking by including questions in traditional and simplified Chinese, and English (Jin et al., 2020). The Massive Multitask Language Understanding (MMLU) benchmark uses the US Medical Licensing Exam (MLE) (Hendrycks et al., 2021) for a subset of tasks. MedMCQA extends these benchmarks to a multilingual evaluation framework (Pal et al., 2022).
@@
Despite the growing interest in Arabic NLP, Arabic medical benchmarks remain limited. Achiam et al. (2023) translated MMLU into 14 languages, including Arabic, with the help of professional human translators. AraSTEM focuses on the task of question-answering and has a medical subset (Mustapha et al., 2024). AraMed similarly presents an Arabic medical corpus and annotated Arabic question-answering dataset (Alasmari et al., 2024). Other datasets are also largely skewed towards the task of Arabic medical question-answering and often have other shortcomings, summarized in Table A1. While these resources are helpful, they do not comprehensively cover the spectrum of Arabic medical tasks, highlighting the need for dedicated benchmarking efforts.
@@
In terms of evaluation, a number of frameworks have been proposed for evaluating clinical AI models. The ‘Governance Model for AI in Healthcare’ consists of four main pillars: fairness, transparency, trustworthiness, and accountability (Reddy et al., 2020). Similarly, Dada et al. (2024) propose the Clinical Language Understanding Evaluation framework, which was designed to assess LLMs with real patient data for various modalities, including patient-specific question-answering, hypothesis deduction, and problem and inquiry summarization. Kanithi et al. (2024) introduce ‘MEDIC’, a framework for evaluating LLMs across five clinically relevant categories: medical reasoning, ethical and bias concerns, data and language understanding, in-context learning, and clinical safety and risk assessment. Although there has been significant progress in the field of benchmarking LLMs in medical tasks, most of the evaluations have been conducted for data in the English language, and no single benchmark exists to assess LLMs in Arabic on more than one medical task and performance metric. There are over 380 million native Arabic speakers (Eberhard et al., 2024), a number of which are monolingual. With the vast potential of LLMs in healthcare, it is crucial to accommodate Arabic-speaking patients to ensure fair deployment.
3 Methodology
@@
Here we describe the details of our methodological framework to construct the datasets and evaluate state-of-the-art LLMs. An overview of MedArabiQ is provided in Figure 1.
3.1 Tasks and Datasets
@@
In order to develop a reliable framework for evaluating LLMs in Arabic healthcare applications, we focused on telehealth consultation and question-answering as the main use cases. These involve not only medical reasoning ability but also natural patient-doctor dialogue. The LLM should mimic the role of a clinician as closely as possible, which includes possession of medical knowledge and the ability to utilize the knowledge, personalized to the needs and queries of patients. Personalization, however, should not lead to bias or prejudice in the LLM’s response, based on the patient’s profile. In fact, the model should show resistance to biased prompts, responding with fairness and objectivity. We derived our datasets from two primary sources: past exams and notes from Arabic medical schools, and the AraMed Dataset (Alasmari et al., 2024). We specifically selected data sources that were unlikely to have been included in prior training datasets.
3.1.1 Multiple Choice Questions
To evaluate the models’ medical understanding, we curated a standard dataset with question-answer pairs, covering foundational and advanced medical topics, such as physiology, anatomy, and neurosurgery. First, we collected paper-based past exams and lecture notes sourced from a large repository of academic materials hosted on student-led social platforms of regional medical schools. No personally identifiable information or real patient data was included, and thus anonymization was not necessary as our data collection complied with privacy and ethical guidelines. These exams and answers were not readily available in structured digital formats, necessitating a rigorous manual process to ensure clarity and correctness. Given that Arabic medical education is not widely digitized, these exams are not publicly accessible in structured formats. Even if some individual questions exist online, the extensive effort required to compile, format, and structure them into a benchmark dataset significantly reduces the likelihood of contamination. Questions were selected to reflect increasing complexity across different academic years, ensuring that model performance could be assessed at varying levels of medical expertise. We selected a random set of 100 multiple-choice questions, digitized and extracted them into CSV files, and performed manual verification. The average question length is 15-30 words.
3.1.2 Multiple Choice Questions with Bias
@@
Following recent work (Schmidgall et al., 2024), we injected bias in multiple choice questions dataset to evaluate how LLMs handle ethical or culturally sensitive scenarios. In particular, we utilized a set of well-defined bias categories proposed by Schmidgall et al. (2024), including (i) confirmation bias, (ii) recency bias, (iii) frequency bias, (iv) cultural bias, (v) false-consensus bias, (vi) status quo bias, and (vii) self-diagnosis bias. By manually injecting the bias, we ensured relevance to the unique linguistic and clinical challenges of Arabic healthcare contexts. This resulted in a dataset consisting of 100 samples.
@@ 3.1.5 Patient-doctor Q&A
AraMed is an Arabic medical corpus for question-answering that was originally sourced from AlTibbi, an online medical patient-doctor discussion forum (Alasmari et al., 2024). The original dataset consists of 400 question-answer pairs. These are all already formulated in the original corpus in Modern Standard Arabic (MSA), ensuring consistency and standardization. We manually selected 100 samples. Our selection process prioritized questions with well-formed queries and meaningful answers, avoiding instances where responses were overly generic (e.g., “Consult a doctor” or “See a specialist”). However, we retained some examples of such cases to reflect real-world user behavior, as patients often seek medical advice even for questions that are better suited for in-person consultations.
Additionally, we maintained proportionality in representation to mirror the dataset’s original structure, particularly for categories like reproductive and sexual health. These categories comprised a significant portion of the dataset and are often underrepresented in Arabic-language medical research despite their importance. Given the sensitivity and cultural taboos associated with these topics, we ensured their inclusion to provide a more realistic and balanced evaluation of the model’s ability to handle diverse medical inquiries. This approach ensures the dataset reflects the varied nature of medical concerns while maintaining its relevance to real-world healthcare scenarios. In order to ensure proportional representation, we selected 100 questions, allocating an equal proportion of samples to each specialty: cardiology, obstetrics and gynecology, surgery, pediatrics, neurology, oncology, endocrinology, dentistry, otholarhyntology, public health, dermatology, primary care, pulmonology, and psychology.
We also incorporated information about the patient, specifically age and gender, when known, into the query. When this information was unknown or erroneous (i.e., the age was an illogical number), it was excluded. Some questions came from an acquaintance of the patient, so no information was known about the patient. The information was usually prepended to the question or inserted after the greeting, in the format of “I am a [man/woman] and I am [x] years old”. The inclusion of this information was useful in incorporating personalization and realism into the case scenarios, making our benchmark dataset better suited to evaluate LLM performance in real-world medical scenarios.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated significant promise for various applications in healthcare. However, their effectiveness in the Arabic medical domain remains unexplored due to the lack of high-quality domain-specific datasets and benchmarks. This study introduces MedArabiQ, a new benchmark dataset consisting of seven Arabic medical tasks, covering multiple specialties and including multiple-choice questions, fill-in-the-blank questions, and patient-doctor questions and answers. We first constructed the dataset using past medical exams as well as publicly available datasets. We conducted an extensive evaluation with eight state-of-the-art open-access and proprietary high-resource LLMs, including GPT-4, Deepseek v3, and Gemini 1.5. Our findings highlight the need for the creation of new high-quality benchmarks that span different languages to ensure fair deployment and scalability of LLMs in healthcare. By establishing this benchmark and releasing the dataset, we provide a foundation for future research aimed at evaluating and enhancing the multilingual capabilities of LLMs for the equitable use of generative AI in healthcare.
Data Availability
In this article, we present a new benchmark dataset, MedArabiQ, designed to evaluate the performance of LLMs on Arabic medical tasks. We release our data for reproducibility to enable a fair evaluation of language models in the future: https://github.com/nyuad-cai/MedArabiQ
@@ 1 Introduction
The use of LLMs in healthcare has sparked enthusiasm due to the potential to improve diagnostic processes, clinical decision-making, and overall patient care (Nazi and Peng, 2023; Meng et al., 2024). One particular application of interest is medical education, in which LLMs can generate concise summaries of information and support interactive learning experiences (Ben´ıtez et al., 2023). To this end, several benchmarks have been proposed to assess the capabilities of LLMs in medical knowledge and reasoning. Despite these advances, challenges remain, including ethical concerns, risks of generating biased or harmful content, and variability in performance in different languages and cultural contexts (Yang et al., 2024; Nazi and Peng, 2023).
@@
Existing benchmarks, such as GLUE and MedQA, cater primarily to English, leaving a significant gap in the evaluation of LLMs for Arabic healthcare applications (Qiu et al., 2024). This is due to multiple factors, including the limited availability of high-quality Arabic datasets for clinical applications, coupled with the unique linguistic complexity of Arabic, especially when considering the many different Arabic dialect regions (e.g., Gulf, Maghreb, Egypt, Levant, among others) in addition to Modern Standard Arabic (MSA) (Salameh et al., 2018). Additionally, although various multilingual models include Arabic in their training data, their performance often falls short in clinical contexts due to insufficient domain-specific resources and a lack of appropriate benchmarks (Nazi and Peng, 2023; Gangavarapu, 2024). Addressing these gaps is critical to unlocking the full potential of LLMs for Arabic-speaking patients and providers, ensuring equitable access to AI-driven advancements in healthcare.
@@
To address these gaps, there is a growing need for frameworks that assess LLM performance in clinical tasks specific to Arabic-speaking populations. By developing benchmarks that reflect real-world clinical interactions, we can ensure more reliable and culturally appropriate LLM deployment in multilingual healthcare systems. In this study, we make several key contributions to address these challenges by introducing MedArabiQ (see Figure 1). First, we developed seven benchmark datasets designed to evaluate LLMs in Arabic healthcare applications while addressing Arabic linguistic complexity along with domain-specific challenges that hinder existing models. We focus on critical healthcare medical tasks, including medical question-answering, clinical dialogue, and ethical decision making. Secondly, we analyze the performance of multilingual and Arabic LLMs, including high-resource proprietary and open-access LLMs, highlighting the impact of linguistic coverage and transparency in training data on healthcare applications. We conduct a comprehensive evaluation to assess model performance, in order to provide a robust foundation for advancing AI-driven solutions for Arabic healthcare tasks.
Generalizable Insights about Machine Learning in the Context of Healthcare
Our work provides a strong foundation for advancing LLM evaluation for Arabic medical tasks. While our work focuses on MSA, future efforts could further develop our preprocessing pipeline to explore the use of regional dialects, which are commonly used in clinical communication. MedArabiQ also eliminates the entry barrier for researchers in underrepresented regions by providing annotated data and an evaluation framework. By promoting open data and reproducibility, we believe that this is an important step towards building LLM agents that are more inclusive and globally relevant. Overall, MedArabiQ supports further research in areas like clinical decision support and telehealth, ultimately contributing to the development of AI models that are better suited for broader adoption in global real-world clinical workflows.
@@ 2 Related Work
Several benchmarks have been proposed for the evaluation of LLMs for various medical tasks, predominantly in the English language. For example, Gao et al. (2023) introduce Dr. Bench, a diagnostic reasoning benchmark for clinical NLP emphasizing clinical text understanding, medical knowledge reasoning, and diagnosis generation. The benchmark combines English-language data from various sources but primarily consists of in-hospital clinical notes. To assess LLM performance in medical question-answering tasks, a number of benchmarks incorporated samples from medical board exams, such as MedQA. MedQA advances multilingual benchmarking by including questions in traditional and simplified Chinese and English (Jin et al., 2020). The Massive Multitask Language Understanding (MMLU) benchmark uses the US Medical Licensing Exam (USMLE) (Hendrycks et al., 2021) for a subset of tasks. MedMCQA extends these benchmarks to a multilingual evaluation framework (Pal et al., 2022).
@@
Despite the growing interest in Arabic NLP, Arabic medical benchmarks remain limited. Achiam et al. (2023) translated MMLU into 14 languages, including Arabic, with the help of professional human translators. AraSTEM focuses on the task of question-answering and includes a medical subset (Mustapha et al., 2024). AraMed similarly presents an Arabic medical corpus and an annotated Arabic question-answering dataset (Alasmari et al., 2024). Other datasets are also largely skewed towards the task of Arabic medical question-answering and have other limitations, summarized in Table A1. While these resources are helpful, they do not comprehensively cover the spectrum of Arabic medical tasks, highlighting the need for dedicated benchmarking efforts.
@@
Clinically, a number of frameworks have been proposed for evaluating clinical AI models. The ‘Governance Model for AI in Healthcare’ consists of four main pillars: fairness, transparency, trustworthiness, and accountability (Reddy et al., 2020). Similarly, Dada et al. (2024) propose the Clinical Language Understanding Evaluation framework, which was designed to assess LLMs with real patient data for various modalities, including patient-specific question-answering, hypothesis deduction, and problem and inquiry summarization. Kanithi et al. (2024) introduce ‘MEDIC’, a framework for evaluating LLMs across five clinically relevant categories: medical reasoning, ethical and bias concerns, data and language understanding, in-context learning, and clinical safety and risk assessment. Although there has been significant progress in the field of benchmarking LLMs in medical tasks, most evaluations have been performed on data in English, and there is no single benchmark to assess LLMs in Arabic on more than one medical task and performance metric. There are over 380 million native Arabic speakers (Eberhard et al., 2024), a number of which are monolingual. With the vast potential of LLMs in healthcare, it is crucial to accommodate Arabic-speaking patients to ensure fair deployment. Further details on related works are present in Appendix A.
3 Methods
In this section, we describe the details of our methodological framework to construct the datasets and evaluate state-of-the-art LLMs. We start by explaining our chosen tasks, models and the experimental setup used to evaluate model performance. In Section 4, we further explain our data collection and preprocessing methods. A general overview of MedArabiQ is provided in Figure 1.
@@ 3 Methods
Figure 1: Overview of MedArabiQ. We construct seven new benchmark datasets and evaluate eight state-of-the-art LLMs on them.
3.1 Tasks and Models
@@
In order to develop a reliable framework for evaluating LLMs in Arabic healthcare applications, we focused on telehealth consultation and question-answering as the main use cases. These involve not only medical reasoning ability but also natural patient-doctor dialogue. The LLM should mimic the role of a clinician as closely as possible, which includes possession of medical knowledge and the ability to utilize the knowledge, personalized to the needs and queries of patients. Personalization, however, should not lead to bias or prejudice in the LLM’s response, based on the patient’s profile.
At the same time, it is important to note that benchmarks used to evaluate LLMs often face the challenge of data contamination. Many modern LLMs are trained on large web crawl datasets sourced from the internet, which can inadvertently include benchmark questions. This overlap inflates performance metrics, making such benchmarks unreliable for fair evaluation (Chunyuan et al., 2024; Dong et al., 2024). Addressing this issue is critical to ensure that benchmark results accurately reflect the true capabilities of LLMs, free from biases introduced by prior exposure to evaluation data.
In addition to categorizing models based on the transparency of their training data, this study further classifies models according to their linguistic coverage. Specifically, we distinguish between models that are broadly multilingual and those that are primarily Arabic-centric multilingual. Although many LLMs include training data from numerous languages, exposure to a language does not inherently confer proficiency (Bender et al., 2021). GPT-4, for example, is known to have been trained on a variety of languages (Achiam et al., 2023). By assessing multilingual and Arabic-centric multilingual models separately, we explore how language coverage and specialization influence performance, particularly in the domain of Arabic medical tasks. Our model grouping is shown in Table B1 of Appendix B.
This grouping reflects important distinctions in model transparency and linguistic specialization, but we acknowledge the considerable heterogeneity within each group regarding architecture, scale, training data, and performance. To better account for this, we provide additional analysis in Figure B1 and Table B2 of the appendix, comparing models by training compute (FLOPs) and dataset size. As expected, larger training compute generally correlates with improved performance, particularly for closed models such as GPT-4 and Gemini 1.5 Pro. Training dataset size shows some correlation within open models, though trends are less consistent. These analyses support the rationale for our grouping by offering further context for interpreting performance differences amid diverse model characteristics.
3.2 Experimental Setup
Our framework evaluates LLMs on Arabic medical tasks through seven datasets derived from two sources: (1) private structured medical exams/notes and (2) public patient-doctor dialogues adopted from AraMed (Alasmari et al., 2024), as will be shown in further detail in the following sections.
Evaluation. The evaluation of LLMs in Arabic healthcare applications requires a comprehensive framework that balances technical performance with real-world applicability (Sallam and Mousa, 2024). Our framework assesses medical reasoning, decision-making, and dialogue-based question-answering (Kanithi et al., 2024; Guo et al., 2023). Multiple-choice questions (MCQ) were evaluated on accuracy of the answer, while open-ended questions, such as those in the fill-in-the-blank and patient-doctor Q&A, were assessed using BERTScore to measure semantic alignment (Zhang et al., 2020). We perform zero-shot prompting for all models and tasks.
Rather than applying a uniform temperature across models and tasks, we tuned temperature settings based on empirical testing and guidance from prior work (Du et al., 2025), which recommends task- and model-aware configuration to improve generation quality without introducing evaluation bias. For classification tasks such as MCQ and fill-in-the-blank, we used a low temperature (0.0 for open models, 0.2 for closed models) to promote deterministic, stable outputs. For generative tasks like patient-doctor Q&A, slightly higher temperatures improved coherence and naturalness without degrading factuality. Accordingly, we used 0.4 for open models (e.g., LLaMA 3.1, Qwen 2.5) and 0.2 for closed models (e.g., GPT-4, Gemini 1.5 Pro). We further extended our experiments with a temperature ablation study across representative models in Appendix B.2. For semantic evaluation, we use the XLM-RoBERTa-Large model in BERTScore computation due to its multilingual training, which includes Arabic, making it more appropriate than English-centric or monolingual Arabic models.
Instruction Prompts. Prompt engineering is crucial for evaluating LLMs. In our prompt engineering experiments, we tested both English and Arabic prompts and found that English prompts were generally more effective. However, for Q&A benchmark tasks, Arabic prompts performed better in open-access models. Based on these findings, we used English prompts for all tasks except for the Q&A datasets. The same prompt was used for each model, though it was customized to each category of tasks as shown in Table C of the appendix. This was crucial to ensure the prompt reflected realistic scenarios as closely as possible, tailored to different use cases.
Answer Processing. For MCQ and fill-in-the-blank with choices tasks, models generate both the index of the predicted correct option and the full textual answer. However, some models, particularly open-access ones, exhibit spelling inconsistencies. To ensure accuracy, we evaluated only the first character generated after the phrase “The correct letter is:”, comparing it against valid options. If the character does not match a valid choice, then the response is deemed invalid. For open-ended benchmarks, where answers are not constrained to multiple-choice formats, we evaluate the full model-generated response.
3.3 Bias Assessment and Mitigation
To ensure that LLMs can be effectively deployed in healthcare and provide meaningful contributions to the field, it is critical to address their potential to replicate human biases. Completely eliminating bias from LLMs is nearly impossible, as the datasets used to train these models are inherently shaped by human judgment, which is itself susceptible to bias. Recognizing this challenge, we developed an evaluation framework to systematically assess model resistance to biases injected into MCQ, measure susceptibility, and evaluate mitigation strategies. The framework draws on methodologies outlined in recent work (Schmidgall et al., 2024), while adapting key elements to the context of Arabic medical datasets and healthcare applications. These adaptations include culturally relevant bias categories, prompts designed for clinical scenarios, and additional evaluation metrics to align with real-world needs.
  1. 1.
  2. 2.
    Bias Testing: Models are tested with biased variations of prompts. We compute the change in accuracy to assess impact of bias (see Table LABEL:tab:bias_examples of the appendix for bias injection examples).
  3. 3.
    See Table E1 in the appendix for an illustration of these mitigation strategies in use.
4 Dataset Construction
As previously stated, our framework for evaluating LLMs in Arabic healthcare applications integrates clinically validated knowledge and linguistically diverse sources to construct a benchmark that mirrors real-world patient needs and clinical reasoning challenges. We do so by deriving our datasets from two primary sources: past exams and notes from Arabic medical schools, and the AraMed Dataset (Alasmari et al., 2024).
4.1 Data Selection
In our data selection process, we specifically selected data sources that were unlikely to have been included in prior training datasets.
4.1.1 Multiple-Choice Questions
To evaluate the models’ medical understanding, we curated a standard dataset with question-answer pairs, covering foundational and advanced medical topics, such as physiology, anatomy, and neurosurgery. From our curated dataset, we selected a random set of 100 multiple-choice questions on an ad-hoc basis by prioritizing: (i) broad coverage of specialties, (ii) varying difficulty levels, and (iii) clarity of question and answer. We then digitized and extracted them into CSV files and performed manual verification. The average question length is 24 words.
4.1.2 Multiple-Choice Questions with Bias
@@
Following recent work (Schmidgall et al., 2024), we injected bias in the multiple-choice questions dataset to evaluate how LLMs handle ethical or culturally sensitive scenarios. In particular, we utilized a set of well-defined bias categories (Schmidgall et al., 2024), including (i) confirmation bias, (ii) recency bias, (iii) frequency bias, (iv) cultural bias, (v) false-consensus bias, (vi) status quo bias, and (vii) self-diagnosis bias. By manually injecting the bias, we ensured relevance to the unique linguistic and clinical challenges of Arabic healthcare contexts. This resulted in a dataset consisting of 100 samples.
@@ 4.1.5 Patient-doctor Q&A
AraMed is an Arabic medical corpus for question-answering that was originally sourced from AlTibbi, an online medical patient-doctor discussion forum (Alasmari et al., 2024). The original dataset consists of 270,000 question-answer pairs, of which 400 were made publicly available. We meticulously selected 100 samples covering the entire range of specialties present originally in AraMed, ensuring quality and avoiding redundancy.